Grade 12General Physics

A man is at a distance of x from the bus..when the bus begins to move with a constant acceleration a with what minimum velocity the man should run towards the bus?

Siddhi pawar
8 Years agoGrade 12

1 Answer

Eshan
8 Years ago
Let$$v$$be the velocity of man so that he just catches the bus.

Consider the whole situation from the buses point of view, It sees the man runnig towards the bus with speed$$v$$, but retardind with acceleration$$a$$ and the initial distance between them is $$x$$. If he is able to ‘just’ catch the bus, his final speed must be zero.

From third equation of motion,
$$v^2-u^2=2ax$$
$$\implies 0^2-v^2=2(-a)x$$ $$\implies v=\sqrt{2ax}$$